OpenStudy (anonymous):

Solve for x. 3x+4-12=28 @One098

OpenStudy (one098):

\[3x+4-12=28\]Okay so first, add 12 to both sides. \[3x+4-12-12=28-12\]The two 12's cancel out, and subtract the two numbers on the right. \[3x+4=40\]Subtract 4 from both sides. \[3x+4-4=40-4\]The two 4's cancel out, and subtract the numbers on the right. \[3x=36\]Divide 3 by both sides. \[\frac{ 3x }{ 3 }=\frac{ 36 }{ 3 }\]The two 3's on the left cancel out and just divide on the right. \[x=12\]So x=12. You can plug it in to check the work.

OpenStudy (anonymous):

thank you so much can you help with others

OpenStudy (one098):

You're most welcome and sure, of course.